2.This question has to do with calculating the multiplier. Please answer questions 2a through 2c. 2a. Define and write the formula for the multiplier. 2b. Compute the multiplier if the MPC-0.80. Interpret the multiplier you just calculated. 2c. Suppose President Obamas economic advisor tells him that the MPC is now 0.50 because stock prices have fallen in the stock market making investors less wealthy now than before stock prices fell in the stock market. Calculate the new multiplier. Interpret the new multiplier you just calculated. Did the multiplier get smaller or bigger when the MPC falls to 0.50? Why? What does this do to consumption and investment for the Stock Market and other goods produced and bought in the U.S. economy? Does disposable income fall or rise when MPC falls from 0.8 to 0.5? Explairn Calculate the new multiplier and compare it with the multiplier you just calculated in 2b and then explain your results.
Answer #1

The theory of multiplier occupies an important place in the modern theory of income and employment.

The concept of multiplier was first of all developed by F.A. Kahn in the early 1930s. But Keynes later further refined it. F.A. Kahn developed the concept of multiplier with reference to the increase in employment, direct as well as indirect, as a result of initial increase in investment and employment.

Keynes, however, propounded the concept of multiplier with reference to the increase in total income, direct as well as indirect, as a result of original increase in investment and income.

Therefore, whereas Kahn’s multiplier is known as ’employment multiplier’, Keynes’ multiplier is known as investment or income multiplier. The essence of multiplier is that total increase in income, output or employment is manifold the original increase in investment. For example, if investment equal to Rs. 100 crores is made, then the income will not rise by Rs. 100 crores only but a multiple of it.

If as a result of the investment of Rs. 100 crores, the national income increases by Rs. 300 crores, multiplier is equal to 3. If as a result of investment of Rs. 100 crores, total national income increases by Rs. 400 crores, multiplier is 4.

The multiplier is, therefore, the ratio of increment in income to the increment in investment. If ?I stands for increment in investment and ?Y stands for the resultant increase in income, then multiplier is equal to the ratio of increment in income (?K) to the increment in investment (?I).

Formula of calculating multiplier: k = ?Y/?I, where k stands for multiplier.

Hence,

k= Multiplier

?Y= resultant increase in income as a result of investment

?I= increment in investment made

2b. Compute the multiplier if MPC= 0.80. Interpret the multiplier you just calculated.

Multiplier = 1/(1-MPC)

Therefore,

It is given that MPC = 0.80, so we substitute the MPC value in the Multiplier formula:

Multiplier= 1/ (1-0.80) = 1 / 0.20 = 5

Therefore, the Multiplier = 5

The result can be interpreted as for an investment made of 1$ the increase in the earning towards GDP will 5$.

We can also say that there will be 5 times increase in earning of the total investment amount.

2c. As the new MPC = 0.50

Hence the new multiplier will be:

Multiplier = 1 / (1-0.50) = 1/ 0.50 = 2

Therefor the new multiplier is 2 when MPC is 0.50.

The interpretation of the new multiplier value is that the earnings contribution towards GDP will be double the value of investments made.

The multiplier gets smaller when MPC falls to 0.50.

As the multiplier falls to 2 the purchasing power for individuals decrease and hence there is decrease in consumption also.

In the stock markets the spending is decreased as a result of decrease in multiplier and hence the stock prices decrease.

The disposable income falls as MPC falls to 0.50 form previous 0.80.